OpenStudy (anonymous):

the distance traveled at a constant speed is directly proportional to the time of travel. If Olivia traveled 165 miles in7.5 hours, how many miles will Olivia travel in 9.6 hours at the same constant speed?

OpenStudy (anonymous):

165/7.5=x/9.6 x=165(9.6)/7.5

OpenStudy (anonymous):

find rate. formula for this problem D = R*T where D =165, R =? and T=7.5 basically: 165 = 7.5R R= 22 Solve for the second prob: D = 22(9.6) D= 211.2